# TRAM: A Transverse Relaxation Time–Aware Qubit Mapping Algorithm for NISQ Devices

### v1

@@ I Introduction
1. Community Detection-assisted Quantum Transverse Relaxation Partitioning (CQTP). CQTP constructs a noise-aware abstraction of the hardware by analysing calibration data, two-qubit gate errors, readout errors, and $T_{2}$ coherence times, to identify regions of qubits that are both densely connected and homogeneously coherent. The resulting partitions form stable physical subgraphs that serve as the substrate for logical circuit placement. This step transforms raw calibration snapshots into topological zones that maximize connectivity while minimizing cross-boundary decoherence.
@@ II Background
In this section, we will give a brief introduction to quantum mechanics, quantum circuits, quantum programs, quantum noise models, and superconducting quantum computers.
@@ II.4 Quantum program and quantum circuit
The quantum program is code written in programming languages to describe quantum computing tasks, incorporating quantum operations and classical control logic. For example, in the Qiskit framework, a quantum circuit is constructed using the QuantumCircuit class [35]. In the following circuit representation, qubits are indexed numerically. For instance, qc.h(0) applies a Hadamard gate to the logical qubit with index 0, denoted as $q_{0}$. Similarly, qc.cx(0, 1) applies a CNOT gate between logical qubit $q_{0}$ with index 0 and $q_{1}$ with index 1, where $q_{0}$ serves as the control qubit and $q_{1}$ as the target qubit. A quantum program is a complete computational task that encompasses quantum circuits, while a quantum circuit is the core computational module describing qubit gate operations. A quantum circuit is a sequence of qubit gates acting on logical qubits, which are mapped to physical qubits during execution. Fig. 1 illustrates the quantum circuit of a decomposed Fredkin gate and its corresponding directed acyclic graph (DAG). Within the DAG, nodes correspond to qubit gates, edges indicate causal dependencies, and the element $cx$ represents a CNOT gate. The circuit depth, defined as the length of the longest path in the DAG, reflects the number of parallelizable operation layers and is a key metric for estimating execution time and decoherence on NISQ devices [36].
@@ II.5 Superconducting quantum computer
Qubit $T_{1}(\mu s)$ $T_{2}(\mu s)$
Readout assignment
error
ID error RX error Pauli-X error CZ error
@@ IV.1 Partitioning method for superconducting quantum chips
The reward function $F$ is defined as the gain from merging two communities. In the initial stage, each physical qubit is a separate community. The CQTP algorithm calculates the reward function $F$ for all possible community merges, and each time merges the two communities that maximize the reward into a new community. This merging process continues until the number of qubits in the new community meets the qubit number requirement of the quantum program. Take a quantum program requiring four physical qubits as an example. Fig. 6(a) illustrates the community merging process. In the initial stage, each physical qubit is a community and serves as a leaf node in the hierarchical tree. Each merging operation targets two existing communities rather than individual nodes, and the new community after merging includes all nodes from the original two communities. The first merge combines $Q_{4}$ and $Q_{5}$ in a new community $\{Q_{4},Q_{5}\}$, and each subsequent merge combines the two communities that maximize the reward $F$ into a new community. The merging process continues until the count of qubits in the new community reaches $4$, resulting in the final community $\{Q_{3},Q_{4},Q_{5},Q_{6}\}$. Fig. 6(b) illustrates the the complete process of hierarchical tree of community merging. In the initial phase, each qubit serves as an individual leaf node. Thereafter, communities that maximize the reward $F$ are merged iteratively until all qubits are consolidated into the root node of the hierarchical tree.
@@ IV.2 Initial mapping
Figure 8: Different routing processes of the same quantum operation qc.cx(0,1) under different initial mapping schemes. (a) Under this initial mapping scheme, no additional SWAP gates are required, enabling the direct execution of quantum operation qc.cx(0,1). (b) Under this initial mapping scheme, three additional SWAP gates are required to realize the quantum operation qc.cx(0,1). The blue bidirectional arrows represent SWAP gates.
In this paper, we propose the Time-Weighted Heatmap-Based Initial Mapping (THIM) algorithm. This algorithm fully considers the qubit transverse relaxation times and the timing position of qubit gates. By adding dynamic weight allocation to two-qubit gates, it incorporates the coherence of physical qubits into the initial mapping consideration. Based on the global cost function, it achieves the optimization of the initial mapping. The THIM algorithm consists of five key stages: DAG construction and gate ordering, time-weighted heatmap generation, noise-aware routing cost calculation, greedy mapping establishment, and global cost optimization. Each stage is designed to address specific challenges in quantum compilation.
@@
Stage 2: Time-Weighted Heatmap Construction. We construct an $N\times N$ symmetric heatmap matrix $\mathrm{HM}\in\mathbb{R}^{N\times N}$ (where $N$ is the number of logical qubits) that quantifies the weighted interaction frequency between logical qubit pairs. Each entry $\mathrm{HM}[q_{i},q_{j}]$ accumulates the time-weighted contributions of all two-qubit gates acting on the logical qubit pair $(q_{i},q_{j})$:
@@
where $TG_{ij}$ is the index set of two-qubit gates on $(q_{i},q_{j})$ ordered in reverse topological time, $k=1,\dots,|TG|$. $\omega_{k}$ is the time weight factor for gate $g_{k}$. The time weight is designed to emphasize gates that execute later in the circuit:
@@ IV.3 Time Adaptive Dynamic SWAP Algorithm
Building on this idea, we propose the Time-Adaptive Dynamic SWAP (T-SWAP) algorithm, which augments the NNC heuristic with noise-awareness. In addition to minimizing routing distance, T-SWAP incorporates qubit-dependent attenuation factors that dynamically track cumulative decoherence exposure, thereby penalizing hotspot qubits and promoting more balanced routing. This discourages SWAP hotspotting, favors qubits with stronger noise resilience, and ultimately reduces both cumulative error rates and depth overhead. A formal definition of the cost function and the full T-SWAP algorithm are given in the following.
@@
To anticipate future dependencies, we also define the look-ahead set $E$, which contains the immediate successors of gates in $F$. Its two-qubit subset is denoted
@@
    Routing by SWAP: If no gate in $F_{2q}$ is executable (i.e. $M=\emptyset$), candidate SWAPs are generated by considering neighbors of the mapped physical qubits in the coupling graph. Each candidate is scored using the T-SWAP cost function $H$ (Eq. 13). The SWAP with the minimum cost is selected, the mapping is updated, and the process continues.
@@ V.1 Evaluation metrics
ibm-perth
dnn_n2, deutsch_n2, quantumwalks_n2,
basis_change_n3, fredkin_n3, linearsolver_n3
ibm-guadalupe
basis_trotter_n4, bell_n4, variational_n4, vqe_n4, hs4_n4,
error_correction3_n5, qaoa_n6, dnn_n8, hhl_n7, vqe_ucces_n4,
vqe_ucces_n6, vqe_ucces_n8
ibm-brooklyn
square_root_n18, qft_n26, qv_n32, multiplier_n45, dnn_n16,
add_n28, wstate_n27, is_n34
@@ V.3 Benchmarking
The benchmarks we used are derived from the examples in QSAMBench as shown in the Tab. 2. The specific quantum tasks and their corresponding simulators are shown in the table. To demonstrate the effectiveness of TRAM, the optimization level used in circuit compilation is set to 0. Only the necessary routing processes are performed to ensure that the circuit is executable on the target quantum device, with no optimizations conducted at either the logical or physical level.
@@ V.4 Evaluation results
On ibm-perth, TRAM attains fidelities of $98.16\%$, $98.13\%$, $97.29\%$, $91.17\%$, $94.68\%$, and $97.31\%$ for the four tested tasks, exceeding SABRE by $1.74\%$, $2.81\%$, $3.20\%$, $2.66\%$, $6.21\%$, and $3.84\%$, respectively. Similarly, on ibm-guadalupe, TRAM achieves fidelities of $90.83\%$, $90.92\%$, $97.27\%$, $96.96\%$, $97.18\%$, and $91.19\%$, surpassing SABRE by $6.01\%$, $0.29\%$, $1.82\%$, $3.45\%$, $5.58\%$, and $5.52\%$. Averaged over all benchmarks, TRAM yields a fidelity gain of $3.59\%$, with the maximum improvement reaching $6.21\%$.
Beyond the numerical gains, TRAM exhibits a stable and noise-adaptive partitioning behavior that translates directly into higher circuit fidelity. In dnn_n2, deutsch_n2, quantumwalks_n2, fredkin_n3, linearsolver_n3, vqe_n4, and bell_n4, TRAM preserves both the gate count and circuit depth of SABRE, yet CQTP identifies qubit partitions with lower intrinsic error rates, thereby enhancing overall fidelity. Notably, in dnn_n2, deutsch_n2, and quantumwalks_n2, where no additional SWAP gates are required to mediate interactions between qubits, the $1.74\%$ improvement arises from selecting a more robust subset of qubits with longer $T_{2}$ coherence times and implementing an optimized initial mapping.
@@
[…]
@@ VI Conclusion
To sum up, addressing the hardware constraints faced by quantum algorithms in NISQ devices and the issue that traditional qubit mapping ignores decoherence noise, this paper proposes the TRAM based on heuristic strategies. By integrating real quantum hardware data, this mechanism constructs highly correlated and noise-resistant partitions, generates an optimized initial mapping layout, and dynamically selects SWAP gate insertion positions based on decoherence noise, effectively making up for the shortcomings of traditional algorithms. Benchmark tests on the Qiskit framework simulator show that, compared with the existing SABRE algorithm, the TRAM algorithm has achieved steady improvement in fidelity, the number of qubit gates, and quantum circuit depth, respectively, as shown in Fig. 11. This significantly optimizes the execution efficiency and reliability of quantum programs. The research results provide new ideas and effective approaches for overcoming the physical limitations of NISQ devices, improving the practical application performance of quantum algorithms, and developing execution tools for compilation. Future research will further explore how to combine real-time noise monitoring with adaptive adjustment strategies in more complex quantum hardware environments to continuously optimize the qubit mapping mechanism and promote the practicalization of quantum computing technology.

### v2

@@ I Introduction
1. Community Detection-Assisted Quantum Transverse Relaxation Partitioning (CQTP). CQTP constructs a noise-aware abstraction of the hardware by analysing calibration data, including two-qubit gate errors, readout errors, and $T_{2}$ coherence times, to identify regions of qubits that are both densely connected and homogeneously coherent. The resulting partitions form stable physical subgraphs that serve as the substrate for logical circuit placement. This step transforms raw calibration snapshots into topological zones that maximize connectivity while minimizing cross-boundary decoherence.
@@ II Background
In this section, we will give a brief introduction to quantum mechanics, quantum circuit, quantum program, quantum noise models, and superconducting quantum computer.
@@ II.4 Quantum program and quantum circuit
The quantum program is code written in programming languages to describe quantum computing tasks, incorporating quantum operations and classical control logic. For example, in the Qiskit framework, a quantum circuit is constructed using the QuantumCircuit class [44]. In the following circuit representation, qubits are indexed numerically. For instance, qc.h(0) applies a Hadamard gate to the logical qubit with index 0, denoted as $q_{0}$. Similarly, qc.cx(0, 1) applies a CNOT between logical qubit $q_{0}$ with index 0 and $q_{1}$ with index 1, where $q_{0}$ serves as the control qubit and $q_{1}$ as the target qubit. A quantum program is a complete computational task that encompasses quantum circuits, while a quantum circuit is the core computational module describing qubit gate operations. A quantum circuit is a sequence of qubit gates acting on logical qubits, which are mapped to physical qubits during execution. Fig. 1 illustrates the quantum circuit of a decomposed Fredkin gate and its corresponding directed acyclic graph (DAG). Within the DAG, nodes correspond to qubit gates, edges indicate causal dependencies, and the element $cx$ represents a CNOT gate. The circuit depth, defined as the length of the longest path in the DAG, reflects the number of parallelizable operation layers and is a key metric for estimating execution time and decoherence on NISQ devices [3].
@@ II.5 Superconducting quantum computer
Qubit $T_{1}(\mu s)$ $T_{2}(\mu s)$ Readout assignment error ID error RX error Pauli-X error CZ error
@@ IV.1 Partitioning method for superconducting quantum chips
The reward function $F$ is defined as the gain from merging two communities. In the initial stage, each physical qubit is a separate community. The CQTP algorithm calculates the reward function $F$ for all possible community merges, and each time merges the two communities that maximize the reward into a new community. This merging process continues until the number of qubits in the new community meets the qubit number requirement of the quantum program. Take a quantum program requiring four physical qubits as an example. Fig. 6(a) illustrates the community merging process. In the initial stage, each physical qubit is a community and serves as a leaf node in the hierarchical tree. Each merging operation targets two existing communities rather than individual nodes, and the new community after merging includes all nodes from the original two communities. The first merge combines $Q_{4}$ and $Q_{5}$ in a new community $\{Q_{4},Q_{5}\}$, and each subsequent merge combines the two communities that maximize the reward $F$ into a new community. The merging process continues until the count of qubits in the new community reaches four, resulting in the final community $\{Q_{3},Q_{4},Q_{5},Q_{6}\}$. Fig. 6(b) illustrates the the complete process of hierarchical tree of community merging. In the initial phase, each qubit serves as an individual leaf node. Thereafter, communities that maximize the reward $F$ are merged iteratively until all qubits are consolidated into the root node of the hierarchical tree.
@@ IV.2 Initial mapping
Figure 8: Different routing processes of the same quantum operation qc.cx(0,1) under different initial mapping schemes. (a) Under this initial mapping scheme, no additional SWAP gates are required, enabling the direct execution of quantum operation qc.cx(0,1). (b) Under this initial mapping scheme, three additional SWAP gates are required to realize the quantum operation qc.cx(0,1). The blue bidirectional arrows represent SWAP gate paths.
In this paper, we propose a Time-Weighted Heatmap-Based Initial Mapping (THIM) algorithm. This algorithm fully considers the qubit transverse relaxation times and the timing position of qubit gates. By adding dynamic weight allocation to two-qubit gates, it incorporates the coherence of physical qubits into the initial mapping consideration. Based on the global cost function, it achieves the optimization of the initial mapping. The THIM algorithm consists of five key stages: DAG construction and gate ordering, time-weighted heatmap generation, noise-aware routing cost calculation, greedy mapping establishment, and global cost optimization. Each stage is designed to address specific challenges in quantum compilation.
@@
Stage 2: Time-Weighted Heatmap Construction. We construct an $N\times N$ symmetric heatmap matrix $\mathrm{HM}\in\mathbb{R}^{N\times N}$ (where $N$ is the number of logical qubits) that quantifies the weighted interaction frequency between logical qubit pairs. Each entry $\mathrm{HM}[q_{i},q_{j}]$ accumulates the time-weighted contributions of all two-qubit gates acting on the logical qubit pair $q_{i},q_{j}]$:
@@
where $TG_{ij}$ is the index set of two-qubit gates on $q_{i},q_{j}]$ ordered in reverse topological time, $k=1,\dots,|TG|$. $\omega_{k}$ is the time weight factor for gate $g_{k}$. The time weight is designed to emphasize gates that execute later in the circuit:
@@ IV.3 Time Adaptive Dynamic SWAP Algorithm
Building on this idea, we propose a Time-Adaptive Dynamic SWAP (T-SWAP) algorithm, which augments the NNC heuristic with noise-awareness. In addition to minimizing routing distance, T-SWAP incorporates qubit-dependent attenuation factors that dynamically track cumulative decoherence exposure, thereby penalizing hotspot qubits and promoting more balanced routing. This discourages SWAP hotspotting, favors qubits with stronger noise resilience, and ultimately reduces both cumulative error rates and depth overhead. A formal definition of the cost function and the full T-SWAP algorithm are given in the following.
@@
To anticipate future dependencies, we also define the look-ahead set $E$, which contains the immediate successors of gates in $F$. Its two-qubit subset is denoted as
@@
    Routing by SWAP: If no gate in $F_{2q}$ is executable (i.e. $M=\varnothing$), candidate SWAPs are generated by considering neighbors of the mapped physical qubits in the coupling graph. Each candidate is scored using the T-SWAP cost function $H$ (Eq. 13). The SWAP with the minimum cost is selected, the mapping is updated, and the process continues.
@@ V.1 Evaluation metrics
ibm-perth dnn_n2, deutsch_n2, quantumwalks_n2, basis_change_n3, fredkin_n3, linearsolver_n3
ibm-guadalupe basis_trotter_n4, bell_n4, variational_n4, vqe_n4, hs4_n4, error_correction3_n5, qaoa_n6, dnn_n8, hhl_n7, vqe_ucces_n4, vqe_ucces_n6, vqe_ucces_n8
ibm-brooklyn square_root_n18, qft_n26, qv_n32, multiplier_n45, dnn_n16, add_n28, wstate_n27, is_n34
@@ V.3 Benchmarking
The benchmarks we used are derived from the examples in QASMBench as shown in the Tab. 2. The specific quantum tasks and their corresponding simulators are shown in the table. To demonstrate the effectiveness of TRAM, the optimization level used in circuit compilation is set to 0. Only the necessary routing processes are performed to ensure that the circuit is executable on the target quantum device, with no optimizations conducted at either the logical or physical level.
@@ V.4 Evaluation results
On ibm-perth, TRAM attains fidelities of $98.16\%$, $98.13\%$, $97.29\%$, $91.17\%$, $94.68\%$, and $97.31\%$ for the six tested tasks, exceeding SABRE by $1.74\%$, $2.81\%$, $3.20\%$, $2.66\%$, $6.21\%$, and $3.84\%$, respectively. Similarly, on ibm-guadalupe, TRAM achieves fidelities of $90.83\%$, $90.92\%$, $97.27\%$, $96.96\%$, $97.18\%$, and $91.19\%$, surpassing SABRE by $6.01\%$, $0.29\%$, $1.82\%$, $3.45\%$, $5.58\%$, and $5.52\%$. Averaged over all benchmarks, TRAM yields a fidelity gain of $3.59\%$, with the maximum improvement reaching $6.21\%$.
Beyond the numerical gains, TRAM exhibits a stable and noise-adaptive partitioning behavior that translates directly into higher circuit fidelity. In dnn_n2, deutsch_n2, quantumwalks_n2, fredkin_n3, linearsolver_n3, vqe_n4, and bell_n4, TRAM preserves both the gate count and circuit depth of SABRE, yet CQTP identifies qubit partitions with lower intrinsic error rates, thereby enhancing overall fidelity. Notably, in dnn_n2, deutsch_n2, and quantumwalks_n2, where no additional SWAP gates are required to mediate interactions between qubits, the $1.74\%$, $2.81\%$, and $3.20\%$ improvements arise from selecting a subset of more robust qubits with longer $T_{2}$ coherence times and implementing an optimized initial mapping.
@@
Table 4 reports the reduction in two-qubit gate count and circuit depth achieved by TRAM relative to SABRE under the same compilation settings. On ibm-guadalupe, relative to SABRE, TRAM reduces the total number of qubit gates by $10.83\%$, $15.38\%$, $7.48\%$, $22.63\%$, $2.54\%$, and $15.42\%$ for the six benchmark tasks, and shortens the circuit depth by $8.92\%$, $10.93\%$, $8.94\%$, $22.00\%$, $5.06\%$, and $18.00\%$, respectively. On ibm-brooklyn, the corresponding gate-count reductions are $14.69\%$, $2.31\%, 11$$1.79\%$, $10.97\%$, $12.99\%$, $12.47\%$, $1.54\%$, and $9.80\%$, while the circuit depth decreases by $10.09\%$, $7.19\%$, $10.30\%$, $8.70\%$, $9.56\%$, $9.28\%$, $24.60\%$, and $18.39\%$, respectively. Overall, relative to SABRE, TRAM achieves an average reduction of $11.49\%$ in gate count and $12.28\%$ in circuit depth on these two backends.
@@ VI Conclusion
To sum up, addressing the hardware constraints faced by quantum algorithms in NISQ devices and the issue that traditional qubit mapping ignores decoherence noise, this paper proposes the TRAM based on heuristic strategies. By integrating real quantum hardware data, this mechanism constructs highly correlated and noise-resistant partitions, generates an optimized initial mapping layout, and dynamically selects SWAP gate insertion positions based on decoherence noise, effectively making up for the shortcomings of traditional algorithms. Benchmark tests on the Qiskit framework simulator show that, compared with the existing SABRE algorithm, the TRAM algorithm has achieved steady improvement in fidelity, the number of qubit gates, and quantum circuit depth, as shown in Fig. 11. This significantly optimizes the execution efficiency and reliability of quantum programs. The research results provide new ideas and effective approaches for overcoming the physical limitations of NISQ devices, improving the practical application performance of quantum algorithms, and developing execution tools for compilation. Future research will further explore how to combine real-time noise monitoring with adaptive adjustment strategies in more complex quantum hardware environments to continuously optimize the qubit mapping mechanism and promote the practicalization of quantum computing technology.
@@ Acknowledgement
DATA AVAILABILITY
The data are not publicly available. The data are available from the authors upon reasonable request.